# No One-Model-Fits-All: Uncovering Spatio-Temporal Forecasting Trade-offs with Graph Neural Networks and Foundation Models

### v1

@@ 1. Introduction
Urban IoT systems such as environmental monitoring platforms rely on short-term weather forecasting (2–6h ahead) techniques to support decision making in mobility planning and environmental control. Densely deployed meteorological sensors such as temperature, pressure or humidity in a wireless sensor network offer rich spatial and temporal coverage but raises questions on how sampling frequency and inter-sensor (node) correlations affect downstream forecasting, an aspect often overlooked by edge-side data acquisition methods like adaptive sampling and filtering (Giouroukis et al., 2020). Forecasting approaches range from statistical models (Autoregressive, Recurrent Neural Network-RNN) to modern architectures (Transformers, Spatio-Temporal Graph Neural Networks- STGNN), each with different sensitivities: Transformers excel at high-frequency data, while RNNs may overfit or underfit depending on resolution (Zerveas et al., 2021). STGNNs explicitly model graph structures to capture spatiotemporal dependencies (Chuyuan et al., 2023; Jin et al., 2023). Most recently, Time Series Foundation Models (TSFMs) like Chronos (Ansari et al., 2024), Moirai (Woo et al., 2024), TimesFM (Das et al., 2024) have emerged, leveraging large-scale pretraining to enable strong zero-shot forecasting across diverse domains. Unlike language, time-series data is highly heterogeneous, making forecasting complex in domains like temperature where dynamic, confounding variables are inherent. TSFMs eliminate ad-hoc modeling by providing a single pre-trained model that works across tasks with zero-shot inference or minimal fine-tuning, demonstrating strong single-sensor univariate forecasting (Mulayim et al., 2024), but their effectiveness in multi-sensor spatiotemporal settings remains an open question. This leads to a critical question: How do explicit graph-based structures (STGNNs) compare against pretrained Time Series Foundation Models, for short-term temperature forecasting under different spatio-temporal conditions? To answer this, we compare STGNNs with VAR, RNNs, Transformers, and two state of the art TSFM models —Moirai (Woo et al., 2024), Chronos (Ansari et al., 2024), TimesFM (Das et al., 2024)—using a 25-node IoT meterological dataset from New Mexico (Gupta et al., 2021; Bhurtel and Rawat, 2025). We vary (1): Temporal resolution through downsampling (5-60 minute intervals) to assess model sensitivity to sampling rates, and (2): Number of nodes (8, 16, 25) to examine robustness across spatial densities. This setup enables comparison of graph vs. non-graph models under varied spatiotemporal conditions. Our findings show that STGNNs outperform classical and Transformer based models when spatial correlations are strong or data is sparse, while TSFMs perform competitively in uniform, high-frequency settings with low spatial variability. This highlights the need to align model choice with data topology and sensing strategy.
@@ 2. Related Works
Graph Neural Networks for time-series forecasting. Time-series forecasting progressed from statistical ARIMA to deep learning, with early sequence models such as Long Short-Term Memory (LSTM) (Tandel and Panickar, 2025) and Gated Recurrent Units (GRU) (an RNN variant) (Gao and Ribeiro, 2022a) captured temporal patterns but largely ignored spatial relationships. Spatio-Temporal GNNs (STGNNs) (Wang et al., 2025) model spatial–temporal dependencies by treating sensors as graph nodes, with graphs built from geographic proximity or cross-sensor similarity (modalities/meteorology). Refinements include STGCN (Huang et al., 2023) and Graph WaveNet (Prabowo et al., 2023), and later attention/Transformer-based STGNNs (Zhao et al., 2023). Despite these advances, such models rely heavily on training data and often generalize poorly to unseen settings.
@@
Time-Series Foundation Models for forecasting. Inspired by the transformative success of Large Language Models (LLMs), Time-Series Foundation Models (TSFMs) have emerged as a promising path toward generalizable time-series analysis, demonstrating strong few-shot and zero-shot generalization across domains like cyber-physical systems (Han et al., 2025), building energy analytics (Mulayim et al., 2025, 2024), and human activity recognition (Nguyen et al., 2025). These models, based on adapted transformer architectures, are pre-trained on large, diverse corpora of time-series data empowering them to perform core tasks like forecasting and anomaly detection. Contemporary TSFMs like MOMENT (Goswami et al., 2024) and TimesFM (Das et al., 2024) aim to generalize across datasets and modalities, but remain limited to a fixed set of tasks. Although these models support covariate inputs as exogenous variables, their integration methods such as flattening multivariate sequences through variate embeddings or using similarity-based attention, have failed to show consistent performance gains as the transformer encoder treats these covriates as extra input tokens alongside the past target history to form context representation, thus, failing to capture the inter-variable relationship. This limitation becomes critical in spatio-temporal environments such as urban sensor networks, where data heterogeneity across sensor types, locations, and sampling resolutions creates complex inter-dependencies that their design is ill-equipped to handle. In contrast, newer generation of true multivariate TSFMs, such as Moirai (Woo et al., 2024), are built on the core assumption of the transformer that all channels in a multivariate series are inherently correlated and interdependent, however, their performanc not studied. Prior work has examined TSFMs, STGNNs (Chuyuan et al., 2023; Jin et al., 2023), and sequential models (Zerveas et al., 2021) in isolation, leaving a critical need for a controlled, cross-paradigm comparison. We rigorously quantify the performance of TSFMs against spatio-temporal Graph Neural Networks, revealing their trade-offs under varied sampling rates and spatial node coverage (or node densities). This analysis establishes a critical empirical baseline, providing essential insight into the practical challenges of data heterogeneity and precisely defining the necessity for effective multi-variate integration in future TSFMs.
@@ 3. Overview of Time Series Forecasting Models
We group models into 3 categories: Classical, TSFMs, and STGNNs. A) Classical baseline models: We include baseline models: Vector AutoRegression (VAR)—a multivariate AR model, Gated Recurrent Unit (GRU)—a sequential RNN variant, and Transformer (Vaswani et al., 2017)—which models temporal dependencies using self-attention.
B) Time Series Foundation Models (TSFMs): TSFMs are large pretrained models designed for zero-shot or transfer forecasting. We evaluate three TSFM models: Moirai, TimesFM, and Chronos. Table 3.1 summarizes their key attributes. Moirai is an encoder-only Transformer trained on 27B observations using mixture log-likelihood loss. It supports multivariate input and models cross-input dependencies via attention, using multi-resolution tokenization (at multi-time scales) over fixed-length patches of a sequence. In contrast, TimesFM is a decoder-only Transformer trained on 100B+ time points (including LOTSA dataset) across energy, weather, and traffic domains. It is univariate but incorporates covariates (auxiliary inputs) and models TS patches via residual blocks, creating tokens to forecast. Similarly, Chronos (Ansari et al., 2024) is a univariate, pretrained TSFM that tokenizes continuous values into discrete bins and uses a transformer to predict future sequences autoregressively. Its Chronos-X(Arango et al., 2025) extension incorporates exogenous covariates through cross-attention mechanisms allowing the model to leverage external variables while maintaining strong zero-shot forecasting capabilities. C) Spatio-Temporal Graph Neural Networks (STGNNs): STGNN models forecast time series by capturing temporal patterns and spatial dependencies using a graph structure, where an adjacency matrix encodes node interactions. STGNNs operate in two learning stages: (i) Temporal learning, where each node processes its historical data using sequential neural network models like Gated Recurrent Units (GRU) or Convolutional Neural Network (CNN) to capture local temporal patterns and (ii) Spatial aggregation, where nodes exchange information with neighbors via Graph Convolutions Network (GCN), with edge weights reflecting inter-node relation. Each node updates its state by combining local and neighbor features. We use 2 STGNNs: (a) GRUGCN (Gao and Ribeiro, 2022b)- first processes each node’s input through GRU units to capture temporal patterns, then applies graph convolutions to aggregate spatial information from neighbor nodes. (b) TGCN (Nantian et al., 2023)- applies a two-layer GCN to the raw input features at each time step to model spatial dependencies, and then uses a GRU to capture temporal patterns. Unlike GRUGCN, the GCN in T-GCN operates on input features rather than GRU outputs. This allows earlier incorporation of neighborhood information. The choice between them depends on whether spatial dependencies should inform temporal processing (TGCN) or follow it (GRUGCN).
@@ 4. Methodology
We used the Internet-of-Battlefield Things (IoBT) data set (Gupta et al., 2021), containing meteorological data (temperature, humidity, rainfall, etc.) from 25 sensor towers covering a 10km x 10km area in New Mexico, sampled every 5 minutes over 9 days. The high sampling frequency provides substantial data volume for short-term forecasting evaluation, with each sensor contributing $9\times 24\times 12=2,592$ samples, totaling 64,800 time steps across all sensors. This volume is well-suited for short-term forecasting where models rely on recent temporal contexts (predicting 4H from 8H histories), and the effective sample size is further amplified through sliding window processing, generating thousands of training sequences for hour-ahead prediction. The dataset is split into 5 training, 2 validation, and 2 testing days, consistent with standard practices for short-term forecasting evaluation (Chuyuan et al., 2023; Das et al., 2024). Temperature forecasting performance is reported as MAE and RMSE averaged across nodes. The sensor network layout and spatial node coverage for $K\in\{8,16,25\}$ nodes are provided in Appendix A.
@@ 5. Experimental Results
Table 5.3. Performance evaluation for STGNNs (GRUGCN and TGCN side-by-side)

### v2

# Evaluating Spatio-Temporal Forecasting Trade-offs Between Graph Neural Networks and Foundation Models

@@ 1. Introduction
Urban IoT systems such as environmental monitoring platforms rely on short-term weather forecasting (2–6h ahead) techniques to support decision making in mobility planning and environmental control. Densely deployed meteorological sensors such as temperature, pressure or humidity in a wireless sensor network offer rich spatial and temporal coverage but raises questions on how sampling frequency and inter-sensor (node) correlations affect downstream forecasting, an aspect often overlooked by edge-side data acquisition methods like adaptive sampling and filtering (adaptive). Forecasting approaches range from statistical models (Autoregressive, Recurrent Neural Network-RNN) to modern architectures (Transformers, Spatio-Temporal Graph Neural Networks- STGNN), each with different sensitivities: Transformers excel at high-frequency data, while RNNs may overfit or underfit depending on resolution (kdd). STGNNs explicitly model graph structures to capture spatiotemporal dependencies (stgnn; stgnn2). Most recently, Time Series Foundation Models (TSFMs) like Chronos (ansari2024chronos), Moirai (moirai), TimesFM (timesFM) have emerged, leveraging large-scale pretraining to enable strong zero-shot forecasting across diverse domains. Unlike language, time-series data is highly heterogeneous, making forecasting complex in domains like temperature where dynamic, confounding variables are inherent. TSFMs eliminate ad-hoc modeling by providing a single pre-trained model that works across tasks with zero-shot inference or minimal fine-tuning, demonstrating strong single-sensor univariate forecasting (buildsys24), but their effectiveness in multi-sensor spatiotemporal settings remains an open question. This leads to a critical question: How do explicit graph-based structures (STGNNs) compare against pretrained Time Series Foundation Models, for short-term temperature forecasting under different spatio-temporal conditions? To answer this, we compare STGNNs with VAR, RNNs, Transformers, and two state of the art TSFM models —Moirai (moirai), Chronos (ansari2024chronos), TimesFM (timesFM)—using a 25-node IoT meterological dataset from New Mexico (iobt; iobt2). We vary (1): Temporal resolution through downsampling (5-60 minute intervals) to assess model sensitivity to sampling rates, and (2): Number of nodes (8, 16, 25) to examine robustness across spatial densities. This setup enables comparison of graph vs. non-graph models under varied spatiotemporal conditions. Our findings show that STGNNs outperform classical and Transformer based models when spatial correlations are strong or data is sparse, while TSFMs perform competitively in uniform, high-frequency settings with low spatial variability. This highlights the need to align model choice with data topology and sensing strategy.
@@ 2. Related Works
Graph Neural Networks for time-series forecasting. Time-series forecasting progressed from statistical ARIMA to deep learning, with early sequence models such as Long Short-Term Memory (LSTM) (lstm3) and Gated Recurrent Units (GRU) (an RNN variant) (gru2) captured temporal patterns but largely ignored spatial relationships. Spatio-Temporal GNNs (STGNNs) (weather3) model spatial–temporal dependencies by treating sensors as graph nodes, with graphs built from geographic proximity or cross-sensor similarity (modalities/meteorology). Refinements include STGCN (stgcn) and Graph WaveNet (wavenet), and later attention/Transformer-based STGNNs (attention4). Despite these advances, such models rely heavily on training data and often generalize poorly to unseen settings.
@@
Time-Series Foundation Models for forecasting. Inspired by the transformative success of Large Language Models (LLMs), Time-Series Foundation Models (TSFMs) have emerged as a promising path toward generalizable time-series analysis, demonstrating strong few-shot and zero-shot generalization across domains like cyber-physical systems (tsfm4), building energy analytics (tsfm2; buildsys24), and human activity recognition (tsfm3). These models, based on adapted transformer architectures, are pre-trained on large, diverse corpora of time-series data empowering them to perform core tasks like forecasting and anomaly detection. Contemporary TSFMs like MOMENT (moment) and TimesFM (timesFM) aim to generalize across datasets and modalities, but remain limited to a fixed set of tasks. Although these models support covariate inputs as exogenous variables, their integration methods such as flattening multivariate sequences through variate embeddings or using similarity-based attention, have failed to show consistent performance gains as the transformer encoder treats these covriates as extra input tokens alongside the past target history to form context representation, thus, failing to capture the inter-variable relationship. This limitation becomes critical in spatio-temporal environments such as urban sensor networks, where data heterogeneity across sensor types, locations, and sampling resolutions creates complex inter-dependencies that their design is ill-equipped to handle. In contrast, newer generation of true multivariate TSFMs, such as Moirai (moirai), are built on the core assumption of the transformer that all channels in a multivariate series are inherently correlated and interdependent, however, their performanc not studied. Prior work has examined TSFMs, STGNNs (stgnn; stgnn2), and sequential models (kdd) in isolation, leaving a critical need for a controlled, cross-paradigm comparison. We rigorously quantify the performance of TSFMs against spatio-temporal Graph Neural Networks, revealing their trade-offs under varied sampling rates and spatial node coverage (or node densities). This analysis establishes a critical empirical baseline, providing essential insight into the practical challenges of data heterogeneity and precisely defining the necessity for effective multi-variate integration in future TSFMs.
@@ 3. Overview of Time Series Forecasting Models
We group models into 3 categories: Classical, TSFMs, and STGNNs. A) Classical baseline models: We include baseline models: Vector AutoRegression (VAR)—a multivariate AR model, Gated Recurrent Unit (GRU)—a sequential RNN variant, and Transformer (transformer)—which models temporal dependencies using self-attention.
B) Time Series Foundation Models (TSFMs): TSFMs are large pretrained models designed for zero-shot or transfer forecasting. We evaluate three TSFM models: Moirai, TimesFM, and Chronos. Table 3.1 summarizes their key attributes. Moirai is an encoder-only Transformer trained on 27B observations using mixture log-likelihood loss. It supports multivariate input and models cross-input dependencies via attention, using multi-resolution tokenization (at multi-time scales) over fixed-length patches of a sequence. In contrast, TimesFM is a decoder-only Transformer trained on 100B+ time points (including LOTSA dataset) across energy, weather, and traffic domains. It is univariate but incorporates covariates (auxiliary inputs) and models TS patches via residual blocks, creating tokens to forecast. Similarly, Chronos (ansari2024chronos) is a univariate, pretrained TSFM that tokenizes continuous values into discrete bins and uses a transformer to predict future sequences autoregressively. Its Chronos-X(chronosX) extension incorporates exogenous covariates through cross-attention mechanisms allowing the model to leverage external variables while maintaining strong zero-shot forecasting capabilities. C) Spatio-Temporal Graph Neural Networks (STGNNs): STGNN models forecast time series by capturing temporal patterns and spatial dependencies using a graph structure, where an adjacency matrix encodes node interactions. STGNNs operate in two learning stages: (i) Temporal learning, where each node processes its historical data using sequential neural network models like Gated Recurrent Units (GRU) or Convolutional Neural Network (CNN) to capture local temporal patterns and (ii) Spatial aggregation, where nodes exchange information with neighbors via Graph Convolutions Network (GCN), with edge weights reflecting inter-node relation. Each node updates its state by combining local and neighbor features. We use 2 STGNNs: (a) GRUGCN (GRUGCN)- first processes each node’s input through GRU units to capture temporal patterns, then applies graph convolutions to aggregate spatial information from neighbor nodes. (b) TGCN (TGCN)- applies a two-layer GCN to the raw input features at each time step to model spatial dependencies, and then uses a GRU to capture temporal patterns. Unlike GRUGCN, the GCN in T-GCN operates on input features rather than GRU outputs. This allows earlier incorporation of neighborhood information. The choice between them depends on whether spatial dependencies should inform temporal processing (TGCN) or follow it (GRUGCN).
@@ 4. Methodology
We used the Internet-of-Battlefield Things (IoBT) data set (iobt), containing meteorological data (temperature, humidity, rainfall, etc.) from 25 sensor towers covering a 10km x 10km area in New Mexico, sampled every 5 minutes over 9 days. The high sampling frequency provides substantial data volume for short-term forecasting evaluation, with each sensor contributing $9\times 24\times 12=2,592$ samples, totaling 64,800 time steps across all sensors. This volume is well-suited for short-term forecasting where models rely on recent temporal contexts (predicting 4H from 8H histories), and the effective sample size is further amplified through sliding window processing, generating thousands of training sequences for hour-ahead prediction. The dataset is split into 5 training, 2 validation, and 2 testing days, consistent with standard practices for short-term forecasting evaluation (stgnn; timesFM). Temperature forecasting performance is reported as MAE and RMSE averaged across nodes. The sensor network layout and spatial node coverage for $K\in\{8,16,25\}$ nodes are provided in Appendix A.
@@ 5. Experimental Results
Table 5.3. Performance evaluation for STGNNs (GRUGCN and TGCN)